OpenStudy (meilendurcer):

Sue needs 2 1/3 cups of flour for a batch of cookies. How many complete batches can she make with 10 cups of flour?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

10/2.33= could make 4 complete batches

OpenStudy (skullpatrol):

@arjont let her do it her self

OpenStudy (meilendurcer):

i was gonna say 5

OpenStudy (skullpatrol):

Why 5?

OpenStudy (meilendurcer):

\[\frac{ 5 }{ 6} + \frac{ 2 }{ 3 }\] =

OpenStudy (skullpatrol):

That does not equal 5

OpenStudy (meilendurcer):

no i mean its the next question i have and the reason i thought the above was 5 was because i divided 10 to 2

OpenStudy (skullpatrol):

What about the $$\frac{1}{3}$$

OpenStudy (meilendurcer):

i was stuck so i came here.

OpenStudy (skullpatrol):

Try this

OpenStudy (skullpatrol):